OpenStudy (anonymous):

When an object floats in water, three-fourths of its volume is beneath the surface. What is the density of the object? 333 kg/m3 1000 kg/m3 750 kg/m3 250 kg/m3

OpenStudy (turingtest):

X% volume above water=X%smaller density than that of water

OpenStudy (anonymous):

so its 250

OpenStudy (turingtest):

25% less dense than water is\[d_W(1-0.25)=1000(1-(0.25))=750\]

OpenStudy (anonymous):

sry misread, i calculated the opposite i had to do 1000-250 lol sry

OpenStudy (turingtest):

haha, no apology necessary